countryboy3642:

A cylinder has a radius of 3 feet and a height of 2.1 feet. What is the approximate volume of the cylinder? Use 3.14 for pi. 12.6 ft³ 59.3 ft³ 237.5 ft³ 8550.2 ft³

Shadow:

Basically it's a circle stacked on top of each other. Solve for the area of the circle, then multiply it by the height.

Shadow:

\[A_{circle} = \pi r^2\] \[V_{cylinder} = A_{circle} \times height\]

countryboy3642:

just give me the answer

Zowie:

The volume (V) of the cylinder with height h and radius r is: V = π r² h π = 3.14 r = 3 ft h = 2.1 ft If you plug in the numbers you get: V = 3.14 * 3² * 2.1 V = 59.346 ft³ ≈ 59.3 ft³ So that would give you the answer of: 59.3 ft³